An experimental study was conducted by a researcher to determine if a new time slot has an effecton the performance of pupils in mathematics. Fifteen randomly selected learners participated in thestudy. Toward the end of the investigation, a standardized assessment was conducted. The samplemean was X = 85 and the standard deviation s = 3. In the standardization of the test, the mean was75 and the standard deviation was 10. Based on the evidence at hand, is the new time slot effective?Use a = 0.05.​

1
Expert's answer
2021-06-25T11:43:47-0400

Difference of two independent normal variables

Let "X" have a normal distribution with mean "\mu_X" and variance "\sigma_X^2."

Let "Y" have a normal distribution with mean "\mu_Y" and variance "\sigma_Y^2."

If "X" and "Y"are independent, then "X-Y"will follow a normal distribution with mean "\mu_X-\mu_Y" and

variance "\sigma_X^2+\sigma_Y^2."


"\mu_{X-Y}=85-75=10"

"s_{X-Y}=\sqrt{(3)^2+(10)^2}=\sqrt{109}"

The following null and alternative hypotheses need to be tested:

"H_0:\mu=0"

"H_1:\mu\not=0"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05,"

"df=n-1=15-1=14" degrees of freedom, and the critical value for a two-tailed test is "t_c=2.144787."

The rejection region for this two-tailed test is "R=\{t:|t|>2.144787\}"

The t-statistic is computed as follows:


"t=\dfrac{\mu_{X-Y}-\mu}{s_{X-Y}/\sqrt{n}}=\dfrac{10-0}{\sqrt{109}/\sqrt{15}}\approx3.709645"

Since it is observed that "|t|=3.709645>2.144787=t_c," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 0, at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value for two-tailed "\alpha=0.05, df=14, t=3.709645" is "p=0.002332," and since "p=0.002332<0.05=\alpha," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 0, at the "\alpha=0.05" significance level.

Therefore, there is enough evidence to claim that the new time slot can be effective, at the "\alpha=0.05" significance level.